[Hallc_running] [Revised Logentry] Hall C RC update: Friday, July 5

gaskelld at jlab.org gaskelld at jlab.org
Fri Jul 5 12:40:02 EDT 2019

Logentry Text:
The near term schedule for beam restoration remains the same since yesterday's update: all cryomodules should be ready by the end of the day today and beam restoration is scheduled for the next several shifts. Beam is not expected in Hall C until Sunday morning.

The LH2 target was condensed this morning, and the hall remains in Controlled Access. Now would be a good time for any last minute equipment checks or repairs.

As noted yesterday, target operators should report for all shifts starting today since we now have a cold LH2 target.  Shift cancellations remain the same as yesterday's update: Shift leader and 3rd shift worker shifts are canceled through Saturday DAY shift;  Saturday SWING and Sunday OWL are on standby. Sunday, DAY shift workers should all report.

We will have an RC meeting today at 4:30 pm in the meeting room on the 2nd floor of the Counting House. Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3698852
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190705/e1485386/attachment.html>

More information about the Hallc_running mailing list